OpenStudy (abbles):

Why is sin(2x) = -2 not possible?

OpenStudy (abbles):

I was given this question: Find all solutions to the equation (sinx - cosx)^2 = 3 And the answer was: There are no solutions. You know that (sinx - cosx)^2 = 1 - sin(2x). Therefore, 1 - sin(2x) = 3 This equation has no solutions since the above equation simplifies to sin(2x) = -2, which is impossible.

OpenStudy (abbles):

Why is it impossible?

OpenStudy (karim728):